OpenStudy (anonymous):

2(5)(5)=5(2)(2) true?

myininaya (myininaya):

false

myininaya (myininaya):

2*5*5>5*2*2

OpenStudy (anonymous):

numbers how many 2 and how many 5 you see there

hero (hero):

lol @ saif

OpenStudy (anonymous):

Sorry i mean 5(5)(2)

OpenStudy (saifoo.khan):

HAHA.

OpenStudy (saifoo.khan):

i was just kiddin.

OpenStudy (anonymous):

tralse

myininaya (myininaya):

2(5)(5)=5(5)(2) is true

OpenStudy (anonymous):

Multiplication is commutative. That means that the individual factors can commute (move around) and it is still the same thing.

myininaya (myininaya):

multiplication is commutative

myininaya (myininaya):

oops i'm too late

OpenStudy (anonymous):

JINX

myininaya (myininaya):

a*b=b*a

OpenStudy (anonymous):

THANKS EVERYONE U ARE LIFESAVERS!

myininaya (myininaya):

a*b*c=c*a*b=c*a*b

myininaya (myininaya):

=c*b*a

myininaya (myininaya):

i repeated myself by accident

OpenStudy (anonymous):

HAHAHA O_o

jhonyy9 (jhonyy9):

false because 50 not is equal 20